There are three cars A, B and C with a mass of 2 tons, 1 ton and 4 tons. The car drove in the same lane, but in different directions. Car A is behind Car B, moving to the right, while from the opposite direction Car C moves to the left at 25 km / h. Car A which has a speed of 20 km / h crashes into car B that moves at 5 km / hour, and finally the two cars collide with car C. If after the streak of collision, the three cars converge and move together (post-crash motion), determine how high the speed of the car and in what direction?
1
Expert's answer
2020-06-15T10:34:47-0400

From the conservation of momentum:


"m_Cv_C-m_Av_A-m_Bv_B=(m_C+m_A+m_B)V"

"4(25)-2(20)-5=(4+2+1)V"

"V=7.86\frac{km}{hr}"